# Digging into contrastive learning for robust depth estimation with diffusion models

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXsubmissionid: 2846conference: Proceedings of the 32th ACM International Conference on Multimedia; October 29, 2024 – November 3, 2024; Canada, Ottawabooktitle: Proceedings of the 32th ACM International Conference on Multimedia (MM ’24), October 29, 2024 – November 3, 2024, Canada, Ottawaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reconstruction
@@ 1. Introduction
Figure 1. Comparisons with RMDE method(Wang et al., 2023b; Saunders et al., 2023; Song et al., 2023) on KITTI and WeatherKITTI. The length of the Line represents the performance degradation magnitude under adverse environments. All the methods are trained on the same dataset WeatherKITTI for fairness.
To better understand the concept of the ‘trinity’, we need to revisit the previous efforts aimed at resolving the challenges of Robust Monocular Depth Estimation (RMDE). Generally, these approaches fall into two main categories: contrastive learning-based alignment approaches(Song et al., 2023; Liu et al., 2021; Zhao et al., 2022a; Wang et al., 2023b) and knowledge distillation-based pseudo-supervised methods(Saunders et al., 2023; Gasperini et al., 2023; Mao et al., 2024). Considering that corruption does not affect the physical scene depth, the former approach typically involves the clear image along with its various augmentations into the network, enforcing consistency among different output depths to provide extra supervision. However, a common issue of these methods is that directly maximizing the depth similarity can easily fall into a collapsing solution(Chen and He, 2020). To address this issue, additional guidance is required, which poses a challenge in a self-supervised manner since the photometric consistency cannot provide a perfect convergence anchor.
@@
In short, we urgently need a perfect label that can not only serve as an anchor for guiding optimization in contrastive learning but also eliminate performance bottlenecks in knowledge distillation. This label seems impossible in self-supervised learning but we ingeniously find it in the diffusion process—the sampled noise. In the forward diffusion process, we obtain noisy images by adding a sampled noise that follows a normal distribution to the original image. During the training process, sampled noise is considered as the GT label for supervising noise prediction. When the prediction is accurate, the original image will be accurately restored(Ho et al., 2020; Song et al., 2022). Therefore, we incorporate this noise as supplementary guidance for contrastive learning to establish a ‘trinity’ contrast pattern. Concretely, we maintain the original consistency constraints unchanged, but on this basis, we apply sampled noise constraints to clear/augment input noise predictions, gathering the pair of estimated noise towards a more stable and precise optimum. If the model is completely robust and accurate, then the three types of noise(i.e., clear/augmented/sampled noise) should be identical, that is, the ‘trinity’.
@@ 2.1. Monocular Depth Estimation
MDE aims to predict the accurate depth from a single image, which is an ill-posed problem. Intuitively, this challenge was initially seen as a dense regression task(Eigen et al., 2014; Eigen and Fergus, 2015) based on GT depth labels. Four years later, Fu et al.(Fu et al., 2018) discovered that reconstructing MDE into a depth bins classification task can significantly improve performance. Still about four years later, as the diffusion model showcases its talents in generative task(Ho et al., 2020; Song et al., 2022), detection(Chen et al., 2022), and segmentation(Wang et al., 2023a), DDP(Ji et al., 2023) first reformulates this task as a depth map denoising task, and lead to giant progress again. Followers like Diffusiondepth(Duan et al., 2023), DDVM(Saxena et al., 2023), VPD(Zhao et al., 2023), TAPD(Kondapaneni et al., 2023), EcoDepth(Patni et al., 2024), Marigold(Ke et al., 2024), and MonoDiffusion(Shao et al., 2023) all demonstrate the advantages of this paradigm in various MDE sub-tasks. D4RD also adopts this promising scheme and extends it on robust depth estimation.
@@ 3.1. Preliminaries
Self-Supervised MDE uses an auxiliary image $\textbf{I}_{a}$ from the neighboring frame to constrains the output depth with a view-synthesis training pattern. Denoting the MDE model as $\mathcal{F}:\textbf{I}\rightarrow\textbf{d}{\in}\mathbb{R}^{W\times H}$, we can reach the warped target frame $\textbf{I}_{a\rightarrow t}$ with:
@@ 4.1. Dataset
WeatherKITTI(Wang et al., 2023b) is built on the KITTI dataset and includes 6 kinds of weather augmentation. We train D4RD on this dataset because, compared with other synthetic datasets(Saunders et al., 2023; Mao et al., 2024), it is more realistic in simulating real-world complex scenarios. WeatherKITTI shares the same train/validate/test splits as KITTI. Therefore, for training, we follow Zhou’s split(Zhou et al., 2017) (containing 19,905 training images and 2,212 validation images), and for testing, we adopt Eigen’s spilt(Eigen et al., 2014) with 697 test images.
@@
Dense(Bijelic et al., 2020) is one of the few real snowy datasets and it has approximately 1,500 snowy images in total. We introduce this benchmark to test the snowy scenes. To ensure consistency with the above dataset, we followed (Wang et al., 2023b) and sampled 1 image every 3 to obtain 500 images as test data. To generate accurate depth GT in challenging snowy conditions, we use the DROR(Charron et al., 2018) algorithm to remove inaccurate depth readings from LiDARs affected by snowflakes. In addition, a small portion of distorted areas has been trimmed, resulting in a final resolution of $1880\times 924$. We will explain the reasons for not using the snowy part of CADC dataset in the supplementary materials.
@@ 4.2. Implementation Details
In this paper, D4RD is based on MonoDiffusion(Shao et al., 2023), which is currently the only diffusion-based self-supervised MDE pipeline that combines the diffusion architecture with Lite-Mono(Zhang et al., 2023). To maintain consistency with previous RMDE researches(Saunders et al., 2023; Wang et al., 2023b; Song et al., 2023), we replace Lite-Mono with MonoViT(Zhao et al., 2022b) in our D4RD.
@@ 4.3. Quantitative and Qualitative Comparison
Here, we adopt the same evaluation metrics and errors up to 80m for all comparisons as (Godard et al., 2019). Please refer to the supplementary for more details. The comparative methods include the baseline MonoDiffusion, traditional RMDE models’ baseline MonoViT(Zhao et al., 2022b), previous SoTA RMDE models WeatherDepth(Saunders et al., 2023), EC-Depth(Song et al., 2023), Robust-Depth(Saunders et al., 2023), and their retained models. Besides, (Wang et al., 2023b) and (Wang et al., 2023b) apply their approaches to various baselines, we want to clarify that we choose the one that uses MonoViT(Zhao et al., 2022b). In (Song et al., 2023), they take a two-stage training pipeline and we select the second stage model for the comparison.

### v3

@@ Abstract.
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXsubmissionid: 2846conference: Conference; Some Days; Anywhere on earthprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reconstruction
@@ 1. Introduction
Figure 1. Comparisons with RMDE method(Wang et al., 2024; Saunders et al., 2023; Song et al., 2023) on KITTI and WeatherKITTI. The length of the Line represents the performance degradation magnitude under adverse environments. All the methods are trained on the same dataset WeatherKITTI for fairness.
To better understand the concept of the ‘trinity’, we need to revisit the previous efforts aimed at resolving the challenges of Robust Monocular Depth Estimation (RMDE). Generally, these approaches fall into two main categories: contrastive learning-based alignment approaches(Song et al., 2023; Liu et al., 2021; Zhao et al., 2022a; Wang et al., 2024) and knowledge distillation-based pseudo-supervised methods(Saunders et al., 2023; Gasperini et al., 2023; Mao et al., 2024). Considering that corruption does not affect the physical scene depth, the former approach typically involves the clear image along with its various augmentations into the network, enforcing consistency among different output depths to provide extra supervision. However, a common issue of these methods is that directly maximizing the depth similarity can easily fall into a collapsing solution(Chen and He, 2020). To address this issue, additional guidance is required, which poses a challenge in a self-supervised manner since the photometric consistency cannot provide a perfect convergence anchor.
@@
In short, we urgently need a perfect label that can not only serve as an anchor for guiding optimization in contrastive learning but also eliminate performance bottlenecks in knowledge distillation. This label seems impossible in self-supervised learning but we ingeniously find it in the diffusion process—the sampled noise. In the forward diffusion process, we obtain noisy images by adding a sampled noise that follows a normal distribution to the original image. During the training process, sampled noise is considered as the GT label for supervising noise prediction. When the prediction is accurate, the original image will be accurately restored(Ho et al., 2020; Song et al., 2020). Therefore, we incorporate this noise as supplementary guidance for contrastive learning to establish a ‘trinity’ contrast pattern. Concretely, we maintain the original consistency constraints unchanged, but on this basis, we apply sampled noise constraints to clear/augment input noise predictions, gathering the pair of estimated noise towards a more stable and precise optimum. If the model is completely robust and accurate, then the three types of noise(i.e., clear/augmented/sampled noise) should be identical, that is, the ‘trinity’.
@@ 2.1. Monocular Depth Estimation
MDE aims to predict the accurate depth from a single image, which is an ill-posed problem. Intuitively, this challenge was initially seen as a dense regression task(Eigen et al., 2014; Eigen and Fergus, 2015) based on GT depth labels. Four years later, Fu et al.(Fu et al., 2018) discovered that reconstructing MDE into a depth bins classification task can significantly improve performance. Still about four years later, as the diffusion model showcases its talents in generative task(Ho et al., 2020; Song et al., 2020), detection(Chen et al., 2022), and segmentation(Wang et al., 2023), DDP(Ji et al., 2023) first reformulates this task as a depth map denoising task, and lead to giant progress again. Followers like Diffusiondepth(Duan et al., 2023), DDVM(Saxena et al., 2023), VPD(Zhao et al., 2023), TAPD(Kondapaneni et al., 2023), EcoDepth(Patni et al., 2024), Marigold(Ke et al., 2024), and MonoDiffusion(Shao et al., 2023) all demonstrate the advantages of this paradigm in various MDE sub-tasks. D4RD also adopts this promising scheme and extends it on robust depth estimation.
@@ 3.1. Preliminaries
Self-Supervised MDE uses an auxiliary image $\textbf{I}_{a}$ from the neighboring frame to constrain the output depth with a view-synthesis training pattern. Denoting the MDE model as $\mathcal{F}:\textbf{I}\rightarrow\textbf{d}{\in}\mathbb{R}^{W\times H}$, we can reach the warped target frame $\textbf{I}_{a\rightarrow t}$ with:
@@ 4.1. Dataset
WeatherKITTI(Wang et al., 2024) is built on the KITTI dataset and includes 6 kinds of weather augmentation. We train D4RD on this dataset because, compared with other synthetic datasets(Saunders et al., 2023; Mao et al., 2024), it is more realistic in simulating real-world complex scenarios. WeatherKITTI shares the same train/validate/test splits as KITTI. Therefore, for training, we follow Zhou’s split(Zhou et al., 2017) (containing 19,905 training images and 2,212 validation images), and for testing, we adopt Eigen’s spilt(Eigen et al., 2014) with 697 test images.
@@
Dense(Bijelic et al., 2020) is one of the few real snowy datasets and it has approximately 1,500 snowy images in total. We introduce this benchmark to test the snowy scenes. To ensure consistency with the above dataset, we followed (Wang et al., 2024) and sampled 1 image every 3 to obtain 500 images as test data. To generate accurate depth GT in challenging snowy conditions, we use the DROR(Charron et al., 2018) algorithm to remove inaccurate depth readings from LiDARs affected by snowflakes. In addition, a small portion of distorted areas has been trimmed, resulting in a final resolution of $1880\times 924$. We will explain the reasons for not using the snowy part of CADC dataset in the supplementary materials.
@@ 4.2. Implementation Details
In this paper, D4RD is based on MonoDiffusion(Shao et al., 2023), which is currently the only diffusion-based self-supervised MDE pipeline that combines the diffusion architecture with Lite-Mono(Zhang et al., 2023). To maintain consistency with previous RMDE researches(Saunders et al., 2023; Wang et al., 2024; Song et al., 2023), we replace Lite-Mono with MonoViT(Zhao et al., 2022b) in our D4RD.
@@ 4.3. Quantitative and Qualitative Comparison
Here, we adopt the same evaluation metrics and errors up to 80m for all comparisons as (Godard et al., 2019). Please refer to the supplementary for more details. The comparative methods include the baseline MonoDiffusion, traditional RMDE models’ baseline MonoViT(Zhao et al., 2022b), previous SoTA RMDE models WeatherDepth(Saunders et al., 2023), EC-Depth(Song et al., 2023), Robust-Depth(Saunders et al., 2023), and their retained models. Besides, (Wang et al., 2024) and (Wang et al., 2024) apply their approaches to various baselines, we want to clarify that we choose the one that uses MonoViT(Zhao et al., 2022b). In (Song et al., 2023), they take a two-stage training pipeline and we select the second stage model for the comparison.